# Metropolis-Hastings Sampling for 3D Gaussian Reconstruction

### v1

@@ 4.2.1 Per-pixel importance field
At every densification step, we first choose a working camera subset $\mathcal{C}_{t}=\{c_{t,1},\dots,c_{t,k_{t}}\}$ from the full training set $\mathcal{C}$ to ensure balanced use of all cameras over time and broad viewing-angle coverage. Note that this selected subset $\mathcal{C}_{t}$ is used only for computing the importance field. The subset size is annealed according to
@@
so that early iterations ($\eta_{t}\!\approx 0$) ensure broad coverage across diverse views, whereas we reduce the subset size to focus on specific vantage points requiring finer coverage. We use a round-robin schedule that walks through $\mathcal{C}$ in contiguous blocks of length $k_{t}$, as this allows every viewpoint to contribute equally to the error signal after a few iterations. For each camera $c\in\mathcal{C}_{t}$, we compare the current prediction $\hat{I}^{(c)}$ to the ground truth image $I^{(c)}$ and record per-pixel SSIM and $\mathcal{L}_{1}$ errors. Averaging over the subset yields multi-view aggregate maps of:
@@
where $k$ is the number of selected cameras, and $p$ indicated image pixels. This setup ensures we capture multi-view errors rather than relying on a single viewpoint. We then feed $\mathrm{SSIM}_{\mathrm{agg}}(p)$ and $\mathcal{L}1_{\mathrm{agg}}(p)$ along with the opacity $O(p)$, which is directly obtained by projecting the per-Gaussian opacity onto pixel $p$, into our importance score. After applying robust normalization to each quantity (to mitigate outliers and match dynamic ranges), the three cues are fused through a logistic function:
@@ 4.2.3 Proposal Distribution
where $\sigma_{\mathrm{fine}}<\sigma_{\mathrm{coarse}}$. The smaller perturbation radius focuses on residual high-importance regions that still yield large weights $I(i)$ (and hence a large $P_{\text{pick}}(i)$) after the coarse phase fills the major coverage gaps. This allows our method to refine the geometry further with locally placed Gaussians.
@@ 4.3 Metropolis-Hastings Proposal Acceptance
where $\Delta\mathcal{L}$ is the photometric change and $c(v^{\prime})$ is the voxel occupancy before the insertion. (See Appendix B for details.) However, evaluating $\mathcal{L}$ requires rendering every view in the camera set $\mathcal{C}_{t}$, which is computationally expensive. Thus, we use the tight empirical correlation between the importance map $I(i)$ and the loss reduction, $-\Delta\mathcal{L}\approx I(i)$, as a surrogate. (See Appendix C for details.)
Moreover, the calculation of $q(\Theta\mid\Theta^{\prime})$ in Eq. 16, requires the enumeration of the reverse action, which is deleting the particular newborn Gaussian. As our framework does not generate such move, we absorb the unavailable reverse-proposal density into the voxel factor $D(v^{\prime})=(1+\lambda_{v}c_{\Theta}(v\textquoteright))^{-1}$, which already penalises insertions in crowded cells. Substituting both surrogates into Eq. 16 gives an upper bound $\min\{1,e^{I(i)}D(v^{\prime})\}$, and mapping the exponential through the logistic $\sigma(z)=\!(1+e^{-z})^{-1}$ yields a practical version of the MH rule:
$\rho(i)=\sigma\!\bigl{(}I(i)\bigr{)}\,D(v\textquoteright)$ ’ ) (18)
Following the Metropolis-Hastings algorithm, we draw a uniform random number $u\sim\mathcal{U}(0,1)$ for each candidate Gaussian. The proposal is accepted and added to the current set $\Theta_{t}$ when $u<\rho(i)$, otherwise it is rejected. Because $\rho(i)=\sigma\!\bigl{(}I(i)\bigr{)}D(v\textquoteright)$ ’ ) is large only when the per-Gaussian importance $I(i)$ is high and the voxel factor $D(v\textquoteright)$ ’ ) is close to one, Gaussians that target significant photometric errors in sparsely populated regions are accepted with high probability, whereas candidates in already crowded voxels or low-error areas are mostly discarded.
@@ 4.4 Comparison with 3DGS-MCMC [20]
Figure 1: PSNR over training iterations for our method and 3DGS-MCMC [20], averaged across all scenes. Dotted vertical lines indicate when each method reaches 98 % of its eventual PSNR, highlighting near-optimal convergence speed. While both methods ultimately attain comparable final PSNR, ours converges substantially faster.
@@
In contrast, our method runs a Metropolis-Hastings birth sampler that targets only positions. We draw global proposals straight in image-space voids highlighted by a multi-view error map, then pass each candidate through a Metropolis-Hastings accept–reject test that weighs its predicted photometric gain against a sparsity prior. This decouples exploration (“where to propose”) from retention (“what to keep”), enabling long jumps to uncovered regions and the principled dismissal of redundant splats, which Langevin-noise walks with unconditional acceptance cannot achieve. Moreover, we analytically derive and rigorously prove that the 3DGS densification process can be recast as a Metropolis–Hastings sampler, establishing a principled bridge between 3DGS and our formulation. In principle, 3DGS-MCMC is a local, always-accept SGLD chain, whereas our approach is a global, importance-driven MH chain focused on the scene’s unexplained areas. This leads to our method achieving faster convergence, see Fig. 1 for the comparison.
@@ 5.1 Experimental Settings
Dataset and Metrics.
We perform comprehensive experiments using the same real-world datasets employed in 3DGS. Specifically, we utilize the scene-scale view synthesis dataset from Mip-NeRF360 [1], which includes nine large-scale real-world scenes comprising five outdoor and four indoor environments. Additionally, we chose two scenes from the Tanks and Temples [21] dataset and two from the Deep Blending [15] dataset, focusing on the same scenes used in the original 3DGS. We evaluate each method using the standard evaluation metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS).
Implementation Details.
Our framework is based on the original 3DGS [18] implementation. We also use the differentiable tile rasterizer implementation provided by 3DGS-MCMC [20]. We conducted experiments on an NVIDIA GeForce RTX 3090 GPU. Averaged across all datasets, our method runs at 0.0528 s per iteration, and 26.4 min for 30000 iterations. During densification, a normalized progress value ($\in[0,1]$) linearly transitions parameters from coarse (larger offsets, voxel sizes, and batch sizes) to fine (smaller offsets and voxel sizes) to shift focus from broad to precise refinements. Coarse proposal scales range from 10.0 to 5.0, fine proposal scales from 2.0 to 1.0, and voxel sizes from 0.02 to 0.005. Batch sizes for coarse and fine proposals are set at 4,500 and 16,000.
Comparison with Baselines.
We perform comparisons with methodologies that model large-scale scenes. We compare our approach against state-of-the-art baselines, such as Plenoxels [10], Instant-NGP [35], Mip-NeRF 360 [1], and 3DGS [18]. For the baselines aside from 3DGS [18], we present the original results as reported in 3DGS [18]. Moreover, we include a comparison with 3DGS-MCMC [20], which has a similar motivation and shows state-of-the-art performance.
5.2 Results
Table 1: Quantitative results of our method and state-of-the-art models evaluated on Mip-NeRF 360 [1], Tanks&Temples [21] and Deep Blending [15] datasets. We present the baseline results from 3DGS [18]. For a fair comparison, we also report the results for 3DGS obtained from our experiments (Denoted as 3DGS*) and the results from 3DGS and 3DGS-MCMC trained with the same number of Gaussians as our method (denoted as 3DGS-S and 3DGS-MCMC, respectively). The unit of the number of Gaussians is million. The results are colored as Aybest , Aysecond-best , and Aythird-best . To avoid redundancy, only 3DGS* results are considered instead of 3DGS.
@@ 5.2 Results
Table 2: Ablations on the proposed contributions for our Per-Gaussian Importance Score calculation. We test on the Tanks & Temples [21] - Truck and Deep Blending [15] - Dr Johnson dataset.
Method | Dataset Truck Dr Johnson
L1 SSIM Opacity PSNR $\uparrow$ # of Gaussians PSNR $\uparrow$ # of Gaussians
25.47 756,180 29.47 766,421
25.41 758,453 29.57 762,765
25.39 743,017 29.52 760,415
25.54 756,104 29.64 766,469
Table 3: Ablation on relocation. Results are presented for both 3DGS and 3DGS-S, the latter using the same number of Gaussians as ours.
Mip-NeRF360 [1] - Bonsai T&T [21] - Train
@@
3DGS [18] 32.14 / 0.940 / 0.206 21.89 / 0.814 / 0.207
3DGS-S [18] 31.37 / 0.935 / 0.205 21.80 / 0.791 / 0.245
Ours w/o Reloc. 32.19 / 0.946 / 0.181 22.16 / 0.811 / 0.219
Ours 32.52 / 0.950 / 0.171 22.44 / 0.821 / 0.198
Figure 2: Qualitative results of our method compared to 3DGS and the corresponding ground truth image from held-out test views. As seen in the red bounding box, our method can better capture the fine and coarse details that 3DGS [18] and 3DGS-MCMC [20] miss.
5.3 Ablation Study
Components for the Importance Scores.
We conducted an ablation study on the key Gaussian components to identify the most influential factors in importance score calculation for accurate sampling. The results, shown in Table 2, support the effectiveness of our Importance Score Computation design. Specifically, combining SSIM, L1 loss, and opacity values leads to higher PSNR scores and reduces the required number of Gaussians compared to both the baseline and variants that rely only on photometric losses. This study underscores the need for careful factor selection. The complete ablation results are presented in Appendix I.
Effectiveness of Relocation.
We explore whether relocating Gaussians was effective in Table 3. Our experiments confirm that the relocation strategy is essential for maintaining high representational fidelity. It shows that dynamically reallocating underperforming Gaussians to areas where they contribute more effectively improves performance in both indoor and outdoor scenes. The complete ablation results for relocation are presented in Appendix H.
@@ 6 Conclusion
In this paper, we present an adaptive sampling framework that unifies densification and pruning for 3D Gaussian Splatting via a probabilistic Metropolis-Hastings scheme. The importance-weighted sampler automatically targets visually or geometrically critical regions while avoiding redundancy. We derive a new and concise probabilistic formulation that links proposals to scene importance, ensuring faster convergence without sacrificing accuracy. Our experiments show that our method matches or surpasses state-of-the-art quality with fewer Gaussians, and our method’s simplicity invites further advances in 3D scene representation.
Limitations and Future Work.
Our method is not without limitations. Our Metropolis-Hastings sampler, though aware of low-density regions, still prioritizes high-density areas, leading to outdoor scenes with highly sparse data performing less well than indoor scenes. Moreover, as we use multi-view photometric errors when proposing points, it adds roughly 5 minutes of extra training time per scene. Crafting proposal distributions tuned for sparse regions and introducing faster optimization to cut the training time would be an exciting direction for future research.
@@
[…]
@@
In Eq. 3, we found out that $\ln p(\mathcal{D}\mid\Theta)\approx-\mathcal{L}(\Theta)$. Combining these two together, we can derive:
@@
If we exponentiate $-\mathcal{E}(\Theta)$:
@@
[…]
@@
Moreover, the sum collapses to just the one term for voxel $v^{\prime}$. Thu,s continuing our derivation:

### v2

@@ 4.2.1 Per-pixel importance field
At every densification step, we first choose a working view subset $\mathcal{C}_{t}=\{c_{t,1},\dots,c_{t,k_{t}}\}$ from the full training set $\mathcal{C}$ to ensure balanced use of all viewpoints over time and broad viewing‐angle coverage. Note that this selected subset $\mathcal{C}_{t}$ is used only for computing the importance field. The subset size is annealed according to
@@
so that early iterations ($\eta_{t}\!\approx 0$) ensure broad coverage across diverse views, whereas we reduce the subset size to focus on specific vantage points requiring finer coverage. We use a round-robin schedule that walks through $\mathcal{C}$ in contiguous blocks of length $k_{t}$, as this allows every viewpoint to contribute equally to the error signal after a few iterations. For each viewpoint $c\in\mathcal{C}_{t}$, we compare the current prediction $\hat{I}^{(c)}$ to the ground truth image $I^{(c)}$ and record per-pixel SSIM and $\mathcal{L}_{1}$ errors. To construct a unified importance field for Gaussian proposal generation, we aggregate these error maps and construct view-averaged importance maps:
@@
where $k$ is the number of selected viewpoints, and $p$ indicated image pixels. This aggregation produces a spatial field that captures the difficulty of reconstruction across the current viewing subset. We then feed $\mathrm{SSIM}_{\mathrm{agg}}(p)$ and $\mathcal{L}1_{\mathrm{agg}}(p)$ along with the opacity $O(p)$, which is directly obtained by projecting the per-Gaussian opacity on pixel $p$, into our importance score. After applying robust normalization to each quantity (to mitigate outliers and match dynamic ranges), the three cues are fused through a logistic function:
@@ 4.2.3 Proposal Distribution
where $\sigma_{\mathrm{fine}}<\sigma_{\mathrm{coarse}}$. The smaller perturbation radius focuses on residual high-importance regions that still yield large weights $I(i)$ (and hence a large $P_{\text{pick}}(i)$) after the coarse phase fills the major coverage gaps. This allows our method to refine the geometry further using locally placed Gaussians. See Fig. 4.4 for a visualization of the coarse–fine proposal phase.
@@ 4.3 Metropolis-Hastings Proposal Acceptance
where $\Delta\mathcal{L}$ is the photometric change and $c(v^{\prime})$ is the voxel occupancy before the insertion. (See Appendix B for details.) However, evaluating $\mathcal{L}$ requires rendering every view in the set $\mathcal{C}_{t}$, which is computationally expensive. Thus, we use the tight empirical correlation between the importance map $I(i)$ and the loss reduction, $\Delta\mathcal{L}\approx I(i)$, as a surrogate. (See Appendix C for details.)
Moreover, the calculation of $q(\Theta\mid\Theta^{\prime})$ in Eq. 18 requires the enumeration of the reverse action, which is deleting the particular newborn Gaussian. As our framework does not generate such a move, we absorb the unavailable reverse-proposal density into the voxel factor $D(v^{\prime})=(1+\lambda_{v}c_{\Theta}(v’))^{-1}$, which already penalises insertions in crowded cells. Substituting both surrogates into Eq. 18 gives an upper bound $\min\{1,e^{I(i)}D(v^{\prime})\}$, and mapping the exponential through the logistic $\sigma(z)=\!(1+e^{-z})^{-1}$ yields a practical version of the MH rule:
[…]
@@ 4.4 Comparison with 3DGS-MCMC [22]
Figure 1: PSNR over training iterations for our method and 3DGS-MCMC, averaged across all scenes. Dotted vertical lines indicate when each method reaches 98 % of its eventual PSNR, highlighting near-optimal convergence speed. While both methods ultimately attain comparable final PSNR, ours converges faster.
Table 1: Time to equal PSNR comparison between our method and 3DGS-MCMC, averaged across all benchmark scenes. Our method consistently reaches target PSNR thresholds faster.
Target PSNR Time - Ours Time - 3DGS-MCMC
(dB) (s / mins) (s / mins)
21 16.30 / 0.27 17.08 / 0.28
24 61.34 / 1.02 98.38 / 1.64
27 287.01 / 4.78 341.64 / 5.69
30 851.52 / 14.19 983.05 / 16.38
Figure 2: 3D visualization of (a) coarse and (b) fine proposal stages (see Sec. 4.2.3 for details).
In contrast, our method runs a Metropolis-Hastings birth sampler that targets only positions. We draw global proposals straight in image-space voids highlighted by a multi-view error map, then pass each candidate through a Metropolis-Hastings accept–reject test that weighs its predicted photometric gain against a sparsity prior. This decouples exploration (“where to propose”) from retention (“what to keep”), enabling long jumps to uncovered regions and the principled dismissal of redundant splats, which Langevin-noise walks with unconditional acceptance cannot achieve. Moreover, we analytically derive and rigorously prove that the 3DGS densification process can be recast as a Metropolis–Hastings sampler, establishing a principled bridge between 3DGS and our formulation.
In principle, 3DGS-MCMC is a local, always-accept SGLD chain, whereas our approach is a global, importance-driven MH chain focused on the scene’s unexplored areas. This leads to our method converging faster, as shown in Fig. 4.4. Measuring time to equal PSNR across all benchmark scenes, our method reaches 30 dB approximately 2.2 minutes faster than 3DGS-MCMC, with consistently faster convergence at intermediate quality thresholds, as shown in Tab. 4.4.
@@ 5.1 Experimental Settings
Dataset and Metrics. We perform comprehensive experiments using the same real-world datasets employed in 3DGS. Specifically, we utilize the scene-scale view synthesis dataset from Mip-NeRF360 [1], which includes nine large-scale real-world scenes comprising five outdoor and four indoor environments. Additionally, we selected two scenes each from the Tanks and Temples [23] and the Deep Blending [17] datasets, using the same scenes as in the original 3DGS. We evaluate each method using the standard evaluation metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS).
Implementation Details. Our framework is based on 3DGS [20] implementation. We also use the differentiable tile rasterizer implementation provided by 3DGS-MCMC [22]. We conducted experiments on an NVIDIA RTX 3090 GPU. Averaged across all datasets, our method runs at 0.0528 seconds per iteration and 26.4 minutes for 30K iterations. During densification, a normalized progress value ($\in[0,1]$) linearly transitions parameters from coarse (larger offsets, voxel sizes, and batch sizes) to fine (smaller offsets and voxel sizes) to shift focus from broad to precise refinements. Coarse proposal scales range from 10.0 to 5.0, fine proposal scales from 2.0 to 1.0, and voxel sizes from 0.02 to 0.005. Batch sizes for coarse and fine proposals are set to 4,500 and 16,000, respectively.
Comparison with Baselines. We compare against methodologies that model large-scale scenes. We compare our approach against state-of-the-art baselines, such as Plenoxels [10], Instant-NGP [37], Mip-NeRF 360 [1], and 3DGS [20]. For the baselines aside from 3DGS, we present the original results as reported in 3DGS. Moreover, we include a comparison with 3DGS-MCMC [22], which has a similar motivation and shows state-of-the-art performance.
Table 2: Quantitative results of our method and state-of-the-art models evaluated on Mip-NeRF 360 [1], Tanks&Temples [23] and Deep Blending [17] datasets. We present the baseline results from 3DGS [20]. For a fair comparison, we also report the results for 3DGS obtained from our experiments (Denoted as 3DGS*), as well as those for 3DGS and 3DGS-MCMC trained with the same number of Gaussians as our method (denoted as 3DGS-S and 3DGS-MCMC, respectively). The unit of the number of Gaussians is million. The results are colored as Aybest , Aysecond-best , and Aythird-best . To avoid redundancy, only 3DGS* results are considered instead of 3DGS.
@@
5.2 Results
Figure 3: Qualitative results of our method compared to 3DGS and the corresponding ground truth image from held-out test views. As shown in the red bounding box, our method better captures the fine and coarse details that 3DGS [20] and 3DGS-MCMC [22] miss.
Table 2 presents the quantitative evaluations, and Fig. 3 demonstrates the qualitative evaluations on real-world scenes. Our method consistently surpasses 3DGS performance with the same number of Gaussians and achieves or exceeds 3DGS performance with fewer points. Unlike 3DGS, which relies on fixed thresholds for densification, our framework leverages a streamlined point-based sampling strategy that drastically reduces storage requirements while preserving high-fidelity reconstructions. In particular, Tab. C in Appendix F demonstrates our approach’s consistent efficiency across varied scene complexities. We also provide the full per-scene evaluations in Appendix E.
Although 3DGS-MCMC can yield slightly higher metrics, it updates and optimizes every Gaussian parameter, which increases its inefficiency and necessitates knowing the scene’s complexity (the number of Gaussians) in advance. In contrast, our method focuses exclusively on sampling and optimizing point positions, providing a simpler yet equally accurate alternative. As evidenced by Fig. 4.4, our method reaches the 98 $\%$ PSNR threshold faster than 3DGS-MCMC, both in iteration and wall-clock time, demonstrating faster convergence. This shows that our framework can still achieve state-of-the-art reconstruction quality even without exhaustive parameter updates.
5.3 Ablation Study
Components for the Importance Scores. We conducted an ablation study on the key Gaussian components to identify the most influential factors in importance score calculation for accurate sampling. The results, shown in Tab. 3, support the effectiveness of our Importance Score Computation design. Specifically, combining SSIM, L1 loss, and opacity values yields higher PSNR scores and reduces the required number of Gaussians compared to both the baseline and variants that rely solely on photometric losses. This study underscores the need for careful factor selection.
Table 3: Ablations on the importance score components.
Method | Dataset Mip-NeRF360 [1] Tanks&Temples [23] Deep Blending [17]
@@ 5.3 Ablation Study
27.22 0.796 0.243 23.89 0.851 0.168 29.87 0.907 0.247
27.30 0.796 0.242 23.92 0.851 0.169 29.85 0.907 0.246
27.30 0.796 0.243 23.87 0.852 0.167 30.02 0.908 0.247
27.34 0.798 0.241 23.99 0.852 0.166 30.12 0.909 0.245
Table 4: Ablations on the core sampling strategy: Metropolis–Hastings (MH) vs. Stochastic Gradient Langevin Dynamics (SGLD). We denote the number of Gaussians next to the method.
Mip-NeRF360 Tanks&Temples Deep Blending
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS + MH (0.85M) 26.87 0.782 0.260 23.66 0.843 0.173 29.46 0.898 0.258
3DGS + SGLD (0.85M) 26.80 0.789 0.250 23.47 0.841 0.178 29.45 0.898 0.259
Table 5: Ablations on relocation and the multi-view importance score. Results presented for 3DGS use the same number of Gaussians as ours.
Mip-NeRF360 Tanks&Temples Deep Blending
@@
3DGS 26.62 0.761 0.299 23.38 0.828 0.214 29.27 0.897 0.272
Ours w/o Relocation 26.94 0.778 0.267 23.69 0.844 0.184 29.57 0.901 0.267
Ours w/o Importance Score 27.31 0.800 0.234 23.79 0.851 0.167 29.97 0.906 0.245
Ours 27.34 0.798 0.241 23.99 0.852 0.166 30.12 0.909 0.245
Effectiveness of Relocation. We explore whether relocating Gaussians was effective in Tab. 5. Our experiments confirm that the relocation strategy is essential for maintaining high representational fidelity. It shows that dynamically reallocating underperforming Gaussians to areas where they contribute more effectively improves performance in both indoor and outdoor scenes.
Contribution of Metropolis-Hastings Sampling. To measure the exact contribution of our core Metropolis-Hastings (MH) sampling strategy, we conducted an ablation study comparing our method directly against the Stochastic Gradient Langevin Dynamics (SGLD) approach from 3DGS-MCMC [22]. We implemented only the core sampling algorithms from each method while maintaining matched Gaussian counts to ensure a fair comparison. The results shown in Tab. 4 demonstrate that MH sampling consistently outperforms SGLD across all benchmark datasets, confirming the effectiveness of our sampling strategy.
Furthermore, to understand the source of our method’s effectiveness, we conducted an ablation study in Tab. 5 that isolates the contributions of the sampling process and the multi-view importance score. Our experiments demonstrate that the method retains strong performance even when the importance scores are removed. This indicates that the underlying sampling mechanism is the primary driver of our approach, while the multi-view importance scores act as an intelligent guidance mechanism.
@@ 6 Conclusion
We present an adaptive sampling framework that unifies densification and pruning for 3D Gaussian Splatting via a probabilistic Metropolis-Hastings scheme. The importance-weighted sampler automatically targets visually or geometrically critical regions while avoiding redundancy. We derive a new, concise probabilistic formulation that links proposals to scene importance, enabling faster convergence without sacrificing accuracy. Our experiments show that our method matches or surpasses state-of-the-art quality with fewer Gaussians, and our method’s simplicity invites further advances in 3D scene representation.
Limitations and Future Work. Our Metropolis-Hastings sampler, though aware of low-density regions, still prioritizes high-density areas, leading to outdoor scenes with sparser data performing worse than indoor scenes. Moreover, using multi-view photometric errors when proposing points adds roughly 5 minutes of extra training time per scene. Crafting multi-scale representations or dual-stream foreground/background samplers, and introducing faster optimization to reduce training time, would be promising directions for future research.
Acknowledgments
This work was supported by IITP grant (RS-2021-II211343: AI Graduate School Program at Seoul National Univ. (5%), RS-2023-00227993: Detailed 3D reconstruction for urban areas from unstructured images (35%), and RS-2025-02303703: Realworld multi-space fusion and 6DoF free-viewpoint immersive visualization for extended reality (60%)) funded by the Korea government (MSIT).
@@ Acknowledgments
$p(\Theta\mid\mathcal{D})\;=\;\frac{p(\mathcal{D}\mid\Theta)\,p(\Theta)}{\underbrace{\int p(\mathcal{D}\mid\Theta)\,p(\Theta)\,d\Theta}_{\displaystyle Z}},$ (21)
@@
In Eq. 5, we found out that $\ln p(\mathcal{D}\mid\Theta)\approx-\mathcal{L}(\Theta)$. Combining these two, we can derive:
@@
If we exponentiate $\mathcal{E}(\Theta)$:
@@
we can derive the unnormalized posterior density. To turn this into a proper probability distribution, we normalize by $Z\;=\;\int e^{-\mathcal{E}(\Theta)}\,d\Theta$, yielding:
@@
Moreover, the sum collapses to a single term for voxel $v^{\prime}$. Thus, continuing our derivation: